# Diversity of rocky planet atmospheres in the C-H-O-N-S-Cl system with interior dissolution, non-ideality, and condensation: Application to TRAPPIST-1e and sub-Neptunes

### v1

@@ Abstract
A quantitative understanding of the nature and composition of low-mass rocky exo(planet) atmospheres during their evolution is needed to interpret observations. The magma ocean stage of terrestrial- and sub-Neptune planets permits mass exchange between their interiors and atmospheres, during which the mass and speciation of the atmosphere is dictated by the planet’s volatile budget, chemical equilibria, and gas/fluid solubility in molten rock. As the atmosphere cools, it is modified by gas-phase reactions and condensation. We combine these processes into an open-source Python package built using JAX called Atmodeller, and perform calculations for planet sizes and conditions analogous to TRAPPIST-1e and K2-18b. For TRAPPIST-1e-like planets, our simulations indicate that CO-dominated atmospheres are prevalent during the magma ocean stage, which, upon isochemical cooling, predominantly evolve into CO2-rich atmospheres of a few hundred bar at 280 K. Around 40% of our simulations predict the coexistence of liquid water, graphite, -sulfur, and ammonium chloride—key ingredients for surface habitability. For sub-Neptune gas dwarfs, pressures are sufficiently high (GPa) to deviate the fugacities of gases from ideality, thereby drastically enhancing their solubilities. This buffers the total atmospheric pressure to lower values than for the ideal case. These effects conspire to produce CH4-rich sub-Neptune atmospheres for total pressures exceeding 3.5 GPa, provided H/C is 100 solar and O2 moderately reducing (3 log10 units below the iron-wüstite buffer). Otherwise, molecular hydrogen remains the predominant species at lower total pressures and/or higher H/C. For all planets at high temperature, solubility enriches C/H in the atmosphere relative to the initial composition.
@@ 1 Introduction
Whilst the discovery of exoplanets has proliferated in recent years, a significant challenge lies in unravelling how their bulk interior compositions relate to atmospheric speciation. Based on current demographics studies, the most abundant kinds of planets are those with masses in between that of Earth and Neptune and are divided into ‘super-Earths’, which are rocky planets with radii less than 1.5 R⊕, and ‘sub-Neptunes’, which are planets with radii larger than 2 R⊕ that likely have H2-rich atmospheres (Bean et al., 2021). Characterising planets with possible molten surfaces (magma oceans)—such as close-in rocky and super-Earth planets in addition to sub-Neptunes with thick hydrogen (H2) envelopes—demands a comprehensive treatment of thermodynamics at the melt-atmosphere interface (Chao et al., 2021; Misener et al., 2023; Rigby et al., 2024; Seo et al., 2024; Werlen et al., 2025). This is because the mass of the melt reservoirs always predominates over that contained in the atmosphere, such that the physicochemical state of the liquid (i.e., the chemical potentials of components set by pressure, temperature and bulk composition) controls the near-surface atmospheric speciation thereby setting a lower boundary condition for the atmosphere (Hirschmann, 2012; Sossi et al., 2020). Above this boundary the atmospheric speciation may be modified by dynamics and chemistry, photochemistry, and escape (Catling & Kasting, 2017; Wordsworth & Kreidberg, 2022; Selsis et al., 2023; Krissansen-Totton et al., 2024; Owen & Schlichting, 2024; Rogers et al., 2024; Cherubim et al., 2025; Nicholls et al., 2025). Ultimately, the spectral signal in the uppermost atmosphere is most accessible for interrogation by space-based observatories such as the James Webb Space Telescope (JWST). Hence, adequately modelling physical and chemical processes at the melt-atmosphere interface is foundational to interpret current and upcoming observations.
@@
During planet formation, volatiles delivered in planetary building blocks (e.g., Alexander et al., 2012; Marty, 2012) and by capture of the nebular gas (e.g., Sharp & Olson, 2022; Young et al., 2023) likely equilibrated with the growing planet’s rocky magma ocean, whose capacity to store volatile species varies according to their solubilities. A solubility law for any given volatile species relates its dissolved concentration in magma to its fugacity, which is, in turn, related to partial pressure through a fugacity coefficient. These laws are determined either experimentally or theoretically via ab initio simulations for a set of conditions including total pressure, temperature, oxygen fugacity and melt composition that affect the stable melt and/or gas species. Solubility studies have focused on deriving relations for major atmosphere-forming species (e.g., H2O, H2, CO2, CO, CH4, N2, SO2, Cl2, and the noble gases) in silicate melts (e.g., Dixon et al., 1995; Newcombe et al., 2017; Sossi et al., 2023; Hirschmann et al., 2012; Ardia et al., 2013; Libourel et al., 2003; Boulliung & Wood, 2022; Bernadou et al., 2021; O’Neill & Mavrogenes, 2022; Thomas & Wood, 2021; Davis & Caracas, 2024; Jambon et al., 1986), and were largely conducted at moderate to high-pressure conditions (hundreds of bar to several GPa) since they were originally motivated by the conditions of Earth’s mantle and crust. These conditions, however, are also comparable to the melt-atmosphere interface of large rocky planets and hence these solubility relations are now finding new applications in exoplanet research.
@@
For interface pressures greater than 1 GPa, gases can no longer be reasonably treated as obeying the ideal gas law, depending on temperature and the nature of the gas species (Bridgman, 1924; Duan & Zhang, 2006). Fugacity coefficients quantify how strongly a gas departs from ideal behaviour and are often expressed in terms of temperature (), pressure (), and composition () (e.g., Redlich & Kwong, 1949; Tian & Heng, 2024). Deviations from unity pertain physically to differences in volume at a given pressure and temperature relative to the ideal gas law, which are products of 1) the finite size of gas particles and 2) attractive/repulsive interactions between them (Van der Waals, 1873) and are themselves a function of , and . They can be derived from real gas equations of state (e.g., Saxena & Fei, 1988; Holland & Powell, 1991), thermodynamic tables and charts (e.g., Chase, 1998), or directly from experimental data and measurements (e.g., Presnall, 1969; Jakobsson & Oskarsson, 1994; Frost & Wood, 1997). Non-ideality is of direct relevance to interpreting the formation and evolution of rocky planets, which are mostly rock by mass and gas by volume. Both gas-phase reactions and gas solubilities are regulated by fugacities, which are only equivalent to partial pressures in the ideal limit. Moreover, pressure (not fugacity) connects to volatile mass balances (e.g., Appendix C, Bower et al., 2022). In essence, non-ideality demands that ‘thermodynamic pressure’, relevant to reactions and solubility, and ‘mechanical pressure’, relevant to mass balance, are treated as distinct entities. This may lead to exoplanet demographics that are notably different from those derived by assuming ideal gas behaviour and chemically inert rocky cores (e.g., Kite et al., 2019).
@@ 2.1 Formulation and numerical solution
Atmodeller is written in the Python programming language for seamless integration with existing modelling frameworks and leverages JAX (Bradbury et al., 2018) for high-performance computation. It consists of three subpackages that contain JAX-compliant solubility laws, real gas equations of state, and thermodynamic data, all of which can be accessed independently and used as components of Python-based modelling frameworks if desired (with or without JAX support). Notably, the automatic differentiation feature of JAX is convenient for its ability to propagate derivative information through solubility laws and real gas equations of state, which often present complex functional forms that may themselves require iterative solution during the overall solution procedure. Another feature, automatic vectorisation, enables Atmodeller to easily spawn batch calculations to efficiently explore parameter space in a single vectorised solve. Finally, just-in-time compilation enables a model to be pre-compiled, which facilitates fast Atmodeller calculations as part of a repeating workflow such as a time integration in evolution models. Atmodeller uses Optimistix (Rader et al., 2024) to solve the root-finding problem and additionally leverages Lineax for linear solves (Rader et al., 2023) and Equinox (Kidger & Garcia, 2021). As with all systems of non-linear equations, obtaining a solution requires that applied constraints (fugacity, mass) can be satisfied simultaneously with the given system of equations (mass balance, equilibrium chemistry) and closure conditions (solubility laws, free energies of individual gas species, real gas equations of state). We independently verified the results of Atmodeller by comparison to FactSage 8.2 (Bale et al., 2016) calculations assuming ideal gas behaviour (Appendix B), and the Python package has additional comparisons with condensates as well as FastChem 3.1.1 (Kitzmann et al., 2024). Furthermore, each solubility law and real gas equation of state implemented in Atmodeller includes an associated test that verifies its output against tabulated data or published figures from the original source, ensuring accuracy and reproducibility of the model components.
@@ 2.3 Real gas equations of state
where is the gas constant, temperature, molar volume, , and the compressibility factor . The lower bound of the integration is the standard state pressure (usually 1 bar or 1 atmosphere, depending on the thermodynamic data source), and conventionally defines an ideal state, i.e. . Typically, experimental data or numerical calculations determine or , which are then fit to an empirical or physically motivated function of temperature and pressure, such as the Compensated Redlich-Kwong EOS (CORK, Holland & Powell, 1991) or the Virial EOS (Shi & Saxena, 1992). Because the behaviour of gases under high is related to their critical temperature and pressure, often, the pressure and temperature dependence of volume is expressed relative to these critical quantities (, ); the law of corresponding states (Van der Waals, 1873). This enables regression of experimental data across similarly behaving species and thereby improves the average fit. For known pressure and temperature, fugacity coefficients and fugacities can then be computed by evaluating the integral in Equation 1. Some functional forms have convenient closed-form solutions of the fugacity, whereas others require numerical solution (see Holland & Powell, 1991, for a discussion). It is also beneficial for numerical stability to ensure the EOS are bounded and smooth, so we extrapolate the compressibility factor, , linearly according to a Virial-type expression (Appendix C).
@@ 3.1.1 Context and parameters
There is significant interest in probing the atmospheres of temperate Earth-sized rocky planets, such as the TRAPPIST-1 planets (Gillon et al., 2017), to understand the nature of potentially habitable environments beyond Earth and the solar system. The earliest outgassed atmospheres form during and immediately following planet assembly, whereby one or several epochs of magma ocean formation is thought to have distributed volatile elements readily between the atmosphere, mantle, and, potentially the metallic core (e.g., Hirschmann, 2012; Lebrun et al., 2013; Sossi et al., 2020; Chen & Jacobson, 2022; Gu et al., 2024; Huang et al., 2024). This establishes the initial budget of volatiles in the atmosphere and interior, which may subsequently participate in geological volatile cycling that further modulates their atmospheric abundances (e.g., Sleep & Zahnle, 2001; Krissansen-Totton et al., 2018; Hakim et al., 2021). Here, we apply Atmodeller to a young TRAPPIST-1e analogue as an example of a small rocky planet during its magma ocean phase following iron-nickel-alloy core formation, although the results are applicable broadly for any Earth-sized rocky planet. These results are used to inform subsequent Atmodeller calculations to determine the post-formation atmosphere and surface abundances of TRAPPIST-1e at its present-day equilibrium temperature (Section 3.1.3).
@@
TRAPPIST-1e has a total mass of 4.13 kg, a mantle mass of 2.91 kg (assuming the same core-mantle mass proportion as Earth), and surface radius of 5861 km (0.9 R⊕) (Agol et al., 2021). To compute the properties of the high temperature atmosphere, we assume a surface temperature of 1800 K and consider 14 volatile species that can partition between the atmosphere and a magma ocean depending on their solubility: H2O (Dixon et al., 1995), H2 (Hirschmann et al., 2012), CO (Yoshioka et al., 2019), CO2 (Dixon et al., 1995), CH4 (Ardia et al., 2013), N2 (Libourel et al., 2003), S-bearing species (Boulliung & Wood, 2022, 2023), Cl-bearing species (Thomas & Wood, 2021). Both the EOS of pure species and the mixing between them are assumed to be ideal. This encompasses the majority of the stable, gaseous molecules of the so-called CHONS group of elements that constitute the majority of living matter. Condensed carbon (graphite) is also allowed to exist depending on thermodynamic stability, since this high temperature condensate, if present, can buffer the fugacities of carbon-bearing species at the most reducing conditions.
@@
We do not explicitly model the metallic core as a reservoir for volatiles, but rather calibrate the initial elemental abundances relative to the bulk silicate Earth (BSE), which is best constrained from natural samples (e.g., Palme & O’Neill, 2014). This allows us to model the exchange of volatiles between the silicate mantle and atmosphere without necessitating comparatively uncertain high , partitioning data that governs equilibrium between silicate and metallic phases. Furthermore, Earth’s core and mantle are expected to be out of equilibrium based on a joint consideration of thermodynamic and fluid dynamic principles (Stevenson, 1981). Nevertheless, our approach does not mandate that the metallic core is volatile free, but rather that it is isolated during the period of atmosphere-mantle exchange that we consider. Alternatively, the core could be thought of as volatile free (relatively, compared to the mantle), if the prescribed abundances are to be correlated 1:1 with the bulk composition.
@@
Therefore, we fix the total abundances of N, S, and Cl based on BSE estimates, scaled to the mass of TRAPPIST-1e. The average BSE estimates for N, S, and Cl are 7.007 kg, 9.47 kg, and 1.09 kg, respectively (Sakuraba et al., 2021; Hirschmann, 2016; Kendrick et al., 2017), and, when scaled to the mantle mass of TRAPPIST-1e, they are 4.85 kg, 6.552 kg, and 7.57 kg, respectively. In a Monte Carlo simulation with 10,000 realisations we uniformly sample, in log10 space, the total amount of hydrogen between 0.1 and 10 Earth oceans, C/H (by mass) between 0.1 and 10, and the oxygen fugacity by 5 log10 units either side of the iron-wüstite (IW) buffer (Appendix D). All simulation data are available to download (Bower, 2025).

### v2

# Diversity of low-mass planet atmospheres in the C-H-O-N-S-Cl system with interior dissolution, non-ideality, and condensation: Application to TRAPPIST-1e and sub-Neptunes

@@ Abstract
A quantitative understanding of the nature and composition of low-mass rocky (exo)planet atmospheres during their evolution is needed to interpret observations. The magma ocean stage of terrestrial- and sub-Neptune planets permits mass exchange between their interiors and atmospheres, during which the mass and speciation of the atmosphere is dictated by the planet’s volatile budget, chemical equilibria, and gas/fluid solubility in molten rock. As the atmosphere cools, it is modified by gas-phase reactions and condensation. We combine these processes into an open source Python package built using JAX called Atmodeller, and perform calculations for planet sizes and conditions analogous to TRAPPIST-1e and K2-18b. For TRAPPIST-1e-like planets, our simulations indicate that CO-dominated atmospheres are prevalent during the magma ocean stage, which, upon isochemical cooling, predominantly evolve into CO2-rich atmospheres of a few hundred bar at 280 K. Around 40% of our simulations predict the coexistence of liquid water, graphite, -sulfur, and ammonium chloride—key ingredients for surface habitability. For sub-Neptune gas dwarfs, pressures are sufficiently high (GPa) that gas fugacities deviate from ideality, thereby drastically enhancing solubilities. This buffers the total atmospheric pressure to lower values than for the ideal case. These effects conspire to produce CH4-rich sub-Neptune atmospheres for total pressures exceeding 3.5 GPa, provided H/C is 100 solar and O2 moderately reducing (3 log10 units below the iron-wüstite buffer). Otherwise, molecular hydrogen remains the predominant species at lower total pressures and/or higher H/C. For all planets at high temperature, solubility enriches C/H in the atmosphere relative to the initial composition.
@@ 1 Introduction
Whilst the discovery of exoplanets has proliferated in recent years, a significant challenge lies in unravelling how their bulk interior compositions relate to atmospheric speciation. Based on current demographics studies, the most abundant kinds of planets are those with masses in between that of Earth and Neptune and are divided into ‘super-Earths’, which are rocky planets with radii less than 1.5 R⊕, and ‘sub-Neptunes’, which are planets with radii larger than 2 R⊕ that likely have H2-rich atmospheres (Bean et al., 2021). Characterising planets with possible molten surfaces (magma oceans)—such as close-in rocky and super-Earth planets in addition to sub-Neptunes with thick hydrogen (H2) envelopes—demands a comprehensive treatment of thermodynamics at the melt-atmosphere interface (Chao et al., 2021; Charnoz et al., 2023; Misener et al., 2023; Rigby et al., 2024; Seo et al., 2024; Werlen et al., 2025). This is because the mass of the melt reservoirs always predominates over that contained in the atmosphere, such that the physicochemical state of the liquid (i.e., the chemical potentials of components set by pressure, temperature and bulk composition) controls the near-surface atmospheric speciation thereby setting a lower boundary condition for the atmosphere (Hirschmann, 2012; Sossi et al., 2020). Above this boundary the atmospheric speciation may be modified by dynamics and chemistry, photochemistry, and escape (Catling & Kasting, 2017; Wordsworth & Kreidberg, 2022; Selsis et al., 2023; Krissansen-Totton et al., 2024; Owen & Schlichting, 2024; Rogers et al., 2024; Cherubim et al., 2025; Nicholls et al., 2025). Ultimately, the spectral signal in the uppermost atmosphere is most accessible for interrogation by space-based observatories such as the James Webb Space Telescope (JWST). Hence, adequately modelling physical and chemical processes at the melt-atmosphere interface is foundational to interpret current and upcoming observations.
@@
During planet formation, volatiles delivered in planetary building blocks (e.g., Alexander et al., 2012; Marty, 2012) and by capture of the nebular gas (e.g., Sharp & Olson, 2022; Young et al., 2023) likely equilibrated with the growing planet’s rocky magma ocean, whose capacity to store volatile species varies according to their solubilities. A solubility law for any given volatile species relates its dissolved concentration in magma to its fugacity, which is, in turn, related to partial pressure through a fugacity coefficient. These laws are determined either experimentally or theoretically via ab initio simulations for a set of conditions including total pressure, temperature, oxygen fugacity and melt composition that affect the stable melt and/or gas species. Solubility studies have focused on deriving relations for major atmosphere-forming species (e.g., H2O, H2, CO2, CO, CH4, N2, SO2, Cl2, and the noble gases) in silicate melts (e.g., Dixon et al., 1995; Newcombe et al., 2017; Sossi et al., 2023; Hirschmann et al., 2012; Ardia et al., 2013; Libourel et al., 2003; Boulliung & Wood, 2022; Bernadou et al., 2021; O’Neill & Mavrogenes, 2022; Thomas & Wood, 2021; Davis & Caracas, 2024; Jambon et al., 1986), and were largely conducted at moderate to high-pressure conditions (hundreds of bar to several GPa) since they were originally devised to cover the conditions of Earth’s mantle and crust. These conditions, however, are also comparable to the melt-atmosphere interface of large rocky planets and hence these solubility relations are now finding new applications in exoplanet research.
@@
For interface pressures greater than 1 GPa, gases can no longer be reasonably treated as obeying the ideal gas law, depending on temperature and the nature of the gas species (Bridgman, 1924; Duan & Zhang, 2006). Fugacity coefficients quantify how strongly a gas departs from ideal behaviour and are often expressed in terms of temperature (), pressure (), and composition () (e.g., Redlich & Kwong, 1949; Tian & Heng, 2024). Deviations from unity pertain physically to differences in volume at a given pressure and temperature relative to the ideal gas law, which are products of 1) the finite size of gas particles and 2) attractive/repulsive interactions between them (Van der Waals, 1873) and are themselves a function of , and . They can be derived from real gas equations of state (e.g., Saxena & Fei, 1988; Holland & Powell, 1991), thermodynamic tables and charts (e.g., Chase, 1998), or directly from experimental data and measurements (e.g., Presnall, 1969; Jakobsson & Oskarsson, 1994; Frost & Wood, 1997). Non-ideality is of direct relevance to interpreting the formation and evolution of rocky planets, which are mostly rock by mass and gas by volume. Both gas-phase reactions and gas solubilities are regulated by fugacities, which are only equivalent to partial pressures in the ideal limit. Meanwhile, pressure (not fugacity) is the operative quantity in satisfying volatile mass balance (e.g., Appendix C, Bower et al., 2022). In essence, non-ideality demands that ‘thermodynamic pressure’, relevant to reactions and solubility, and ‘mechanical pressure’, relevant to mass balance, are treated as distinct entities. This may lead to exoplanet demographics that are notably different from those derived by assuming ideal gas behaviour and chemically inert rocky cores (e.g., Kite et al., 2019).
@@ 2.1 Formulation and numerical solution
Atmodeller is written in the Python programming language for seamless integration with existing modelling frameworks and leverages JAX (Bradbury et al., 2018) for high-performance computation. It consists of three subpackages that contain JAX-compliant solubility laws, real gas equations of state, and thermodynamic data, all of which can be accessed independently and used as components of Python-based modelling frameworks if desired (with or without JAX support). Notably, the automatic differentiation feature of JAX is convenient for its ability to propagate derivative information through solubility laws and real gas equations of state, which often present complex functional forms that may themselves require iterative solution during the overall solution procedure. Another feature, automatic vectorisation, enables Atmodeller to easily spawn batch calculations to efficiently explore parameter space in a single vectorised solve. Finally, just-in-time compilation enables a model to be pre-compiled, which facilitates fast Atmodeller calculations as part of a repeating workflow such as a time integration in evolution models. Atmodeller uses Optimistix (Rader et al., 2024) to solve the root-finding problem and additionally leverages Lineax for linear solves (Rader et al., 2023) and Equinox (Kidger & Garcia, 2021). As with all systems of non-linear equations, obtaining a solution requires that applied constraints (fugacity, mass) can be satisfied simultaneously with the given system of equations (mass balance, equilibrium chemistry) and closure conditions (solubility laws, free energies of individual gas species, real gas equations of state). We independently verified the results of Atmodeller by comparison to FactSage 8.2 (Bale et al., 2016) calculations assuming ideal gas behaviour (Appendix C), and the Python package has additional comparisons with condensates as well as FastChem 3.1.1 (Kitzmann et al., 2024). Furthermore, each solubility law and real gas equation of state implemented in Atmodeller includes an associated test that verifies its output against tabulated data or published figures from the original source, ensuring accuracy and reproducibility of the model components.
@@ 2.3 Real gas equations of state
where is the gas constant, temperature, molar volume, , and the compressibility factor . The lower bound of the integration is the standard state pressure (usually 1 bar or 1 atmosphere, depending on the thermodynamic data source), and conventionally defines an ideal state, i.e. . Typically, experimental data or numerical calculations determine or , which are then fit to an empirical or physically motivated function of temperature and pressure, such as the Compensated Redlich-Kwong EOS (CORK, Holland & Powell, 1991) or the Virial EOS (Shi & Saxena, 1992). Because the behaviour of gases under high is related to their critical temperature and pressure, often, the pressure and temperature dependence of volume is expressed relative to these critical quantities (, ); the law of corresponding states (Van der Waals, 1873). This enables regression of experimental data across similarly behaving species and thereby improves the average fit. For known pressure and temperature, fugacity coefficients and fugacities can then be computed by evaluating the integral in Equation 1. Some functional forms have convenient closed-form solutions of the fugacity, whereas others require numerical solution (see Holland & Powell, 1991, for a discussion). It is also beneficial for numerical stability to ensure the EOS are bounded and smooth, so we extrapolate the compressibility factor, , linearly according to a Virial-type expression (Appendix D).
@@ 3.1.1 Context and parameters
There is significant interest in probing the atmospheres of temperate Earth-sized rocky planets, such as the TRAPPIST-1 planets (Gillon et al., 2017), , to understand the nature of potentially habitable environments beyond Earth and the solar system. The earliest outgassed atmospheres form during and immediately following planet assembly, whereby one or several epochs of magma ocean formation is thought to have distributed volatile elements readily between the atmosphere, mantle, and, potentially the metallic core (e.g., Hirschmann, 2012; Lebrun et al., 2013; Sossi et al., 2020; Chen & Jacobson, 2022; Gu et al., 2024; Dasgupta et al., 2024; Huang et al., 2024). This establishes the initial budget of volatiles in the atmosphere and interior, which may subsequently participate in geological volatile cycling that further modulates their atmospheric abundances (e.g., Sleep & Zahnle, 2001; Krissansen-Totton et al., 2018; Hakim et al., 2021). Here, we apply Atmodeller to a young TRAPPIST-1e analogue as an example of a small rocky planet during its magma ocean phase following iron-nickel-alloy core formation, although the results are applicable broadly for any Earth-sized rocky planet. These results are used to inform subsequent Atmodeller calculations to determine the post-formation atmosphere and surface abundances of TRAPPIST-1e at its present-day equilibrium temperature (Section 3.1.3).
@@
TRAPPIST-1e has a total mass of 4.13 kg, a mantle mass of 2.91 kg (assuming the same core-mantle mass proportion as Earth), and surface radius of 5861 km (0.9 R⊕) (Agol et al., 2021). To compute the properties of the high temperature atmosphere, we assume a surface temperature of 1800 K, We consider 14 gas species: This encompasses the majority of the stable, gaseous molecules of the so-called CHONS group of elements, in addition to Cl, that constitute the majority of living matter. .
Of the gas species listed above, the following solubility laws are used to distribute them between the atmosphere and the magma ocean: H2O (Dixon et al., 1995), H2 (Hirschmann et al., 2012), CO (Yoshioka et al., 2019), CO2 (Dixon et al., 1995), CH4 (Ardia et al., 2013), N2 (Libourel et al., 2003), S-bearing species (Boulliung & Wood, 2022, 2023), Cl-bearing species (Thomas & Wood, 2021). Both the EOS of pure species and the mixing between them are assumed to be ideal. Condensed carbon (graphite) is also allowed to exist depending on thermodynamic stability, since this high temperature condensate, if present, can buffer the fugacities of carbon-bearing species at the most reducing conditions.
@@
We do not explicitly model the metallic core as a reservoir for volatiles, but rather calibrate the initial elemental abundances relative to the bulk silicate Earth (BSE), which is best constrained from natural samples (e.g., Palme & O’Neill, 2014). This allows us to model the exchange of volatiles between the silicate mantle and atmosphere without necessitating comparatively uncertain high , partitioning data that governs equilibrium between silicate and metallic phases. Nevertheless, our approach does not mandate that the metallic core is volatile free, but rather that it is isolated during the period of atmosphere-mantle exchange that we consider. Alternatively, the core could be thought of as volatile free (relatively, compared to the mantle), if the prescribed abundances are to be correlated 1:1 with the bulk composition.
@@
Therefore, we fix the total abundances of N, S, and Cl based on BSE estimates, scaled to the mass of TRAPPIST-1e. The average BSE estimates for N, S, and Cl are 7.01 kg, 9.47 kg, and 1.09 kg, respectively (Sakuraba et al., 2021; Hirschmann, 2016; Kendrick et al., 2017), and, when scaled to the mantle mass of TRAPPIST-1e, they are 4.85 kg, 6.55 kg, and 7.57 kg, respectively. In a Monte Carlo simulation with 10,000 realisations we uniformly sample, in log10 space, the total amount of hydrogen between 0.1 and 10 Earth oceans , C/H (by mass) between 0.1 and 10, and the oxygen fugacity by 5 log10 units either side of the iron-wüstite (IW) buffer (Appendix E). All simulation data are available to download (Bower, 2025).